# Improving Zero-Shot Chinese-English Code-Switching ASR with $k$NN-CTC and Gated Monolingual Datastores

### v2

@@ Abstract
1 Introduction
In today’s increasingly interconnected world, the ability to accurately transcribe and understand speech in multilingual and code-switching (CS) environments is of paramount importance [1]. Automatic speech recognition (ASR) systems play a crucial role in facilitating communication across linguistic boundaries. Recent years have witnessed rapid advancements in ASR technology. However, traditional ASR models typically rely on vast amounts of labeled data for each language or dialect, posing a significant challenge in scenarios where such data is scarce or unavailable [2]. Consider, for instance, a scenario where speakers seamlessly transition between two languages, such as English and Chinese, within the same conversation. Traditional ASR systems trained on monolingual datasets struggle to accurately transcribe this CS speech, as similar pronunciations across languages could confuse the model.
@@ 1 Introduction
Addressing this, researchers have explored zero-shot learning to enable model generalization to new tasks or languages without specific training data. Yan et al. [13] suggest a cross-lingual pseudo-labeling modification for monolingual modules to produce transliterations of foreign speech, aiming to circumvent the error propagation of frame-wise language identification (LID) decisions. Despite this advancement, their approach necessitates fine-tuning each monolingual model with cross-lingual pseudo-labeling and using dual encoders during decoding, highlighting ongoing challenges in CS-ASR development.
@@ 2.1 CS-ASR based on $k$NN-CTC
2.2 $k$NN-CTC with gated monolingual datastores
@@
3 Experimental setup
3.1 Dataset
We utilize ASCEND [20], a Chinese-English dataset for CS-ASR. We split the training set of ASCEND into three parts: Chinese, English, and Mixed (CS-ASR data), denoted as $S_{CN}$, $S_{EN}$, $S_{MIX}$ respectively. Our experiments only employ the Chinese and English subsets for fine-tuning the models and constructing datastores to keep the zero-shot setting. Additionally, we utilize the ASCEND test set (denoted as TEST) and the mixed training set ($S_{MIX}$) as our test sets. The details of each subset are shown in Table 1.
Table 1: Details of splitting the dataset.
@@ 3.1 Dataset
Table 2: MER (%) of our proposed method based on the Conformer fine-tuned with $S_{CN}$ and $S_{EN}$.
@@
Table 3: MER (%) of our proposed method based on the Wav2vec2-XLSR fine-tuned with $S_{CN}$ and $S_{EN}$.
@@
3.2 Implementation details
Our experiments are conducted using the open-source toolkit WeNet [21] for Conformer [22] and HuggingFace’s Transformers [23] for Wav2vec2-XLSR [24].
The Conformer-CTC baseline employs log Mel-filter banks, with a 25ms window and 10ms shift, as input. We utilize the open-source checkpoint111 https://github.com/wenet-e2e/wenet/blob/main/docs/pretrained_models.en.md pre-trained by WenetSpeech [25], comprising 12 blocks of Conformer encoder layers with 8 attention heads and 3 blocks of bi-transformer decoder layers. During finetuning, a learning rate of $5\times 10^{-5}$ and a batch size of 16 with 5000 warmup steps on 4 GTX 3090 GPUs are employed, with weights of 0.3 assigned to the CTC loss and 0.7 to the attention loss.
@@ 3.2 Implementation details
In our experimentation with the Wav2vec2-XLSR-CTC baseline, we leverage the open-source checkpoint [^2] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-chinese-zh-cn and fine-tune the baseline with codebase [^3] https://github.com/HLTCHKUST/ASCEND provided by ASCEND [20]. We abbreviate Conformer-CTC to Conformer and Wav2vec2-XLSR-CTC to Wav2vec2-XLSR for simplicity. All reported results are derived using CTC greedy search decoding. It is important to note that our models are fine-tuned exclusively on the subsets $S_{CN}$ and $S_{EN}$, thus preserving the zero-shot CS-ASR setting.
We reimplement the version of $k$NN-CTC (full) using FAISS [26], which has demonstrated superior performance in in-domain ASR tasks [16]. For Conformer, we follow $k$NN-CTC to determine the location of keys by extracting the output of the final encoder layer’s feedforward network (FFN). For Wav2vec2-XLSR, we select the output of the Wav2vec2-XLSR encoder as the location of keys. We set $k$ to 1024 following $k$NN-CTC for both baselines.
Regarding the gated monolingual datastore selection mechanism, we set $n$=300 for the Conformer baseline and $n$=10 for the Wav2vec2-XLSR baseline to compute the average distances of the $n$ nearest neighbors from $D_{CN}$ and $D_{EN}$. We set the distribution calibration temperature $t$ to 5 for the Conformer baseline and 200 for the Wav2vec2-XLSR baseline, adjusting the probability distribution for the alternative language.
We adopt the Mixture Error Rate (MER) as outlined in [27] for metrics. MER accounts for both Chinese characters and English words when calculating the edit distance. Errors are tallied separately for Chinese and English based on the language of the reference token.
4 Results
4.1 Zero-shot CS-ASR results
@@
The evaluation results for the Conformer and Wav2vec2-XLSR baselines are summarized in Table 2 and Table 3, respectively. Our approach, employing the $k$NN-CTC with a single datastore $D_{ALL}$, as well as our method utilizing dual monolingual datastores $D_{CN}$ and $D_{EN}$, outperform the fine-tuned CTC method across both baseline models and test sets. Moreover, our integrated method, incorporating dual monolingual datastores with a gated datastore selection mechanism, demonstrates superior performance in all scenarios. This highlights that the utilization of the bilingual datastore introduces distracting noise during the retrieval process, which adversely affects performance. By implementing the gated mechanism to select a monolingual datastore, we effectively mitigate interference from alternate languages, resulting in precise retrieval and improved performance. Our approach achieves a relative MER reduction of 4.4% and 7.4% on the $TEST$ and $S_{MIX}$ sets respectively for the Conformer baseline. Furthermore, for the Wav2vec2-XLSR baseline, we observe a relative MER reduction of 3.6% and 5.1% on the two test sets separately.
@@
Figure 2: Visualization of average distances $d_{CN}$ and $d_{EN}$. The green dotted vertical line represents occurrences of CS. The color blue represents Chinese, while orange represents English.
@@
4.2 Ablation study
Figure 3 illustrates the results of tuning hyperparameters $n$ and $t$ on the $DEV$ set. Setting $n$ to 1 corresponds to considering only the shortest distance. It is observed that performance improves as $n$ increases, peaking at $n=10$, beyond which performance gradually declines. The scale temperature $t$, as defined in Equation 9, determines the extent to which the probability distribution of the alternative language is suppressed. As $t$ increases, MER is stably decreasing. When $t$ is up to 200 and 500, it achieves the best performance and does not improve anymore, indicating that the selection mechanism is accurate enough to select the corresponding language datastore for each frame.
The results of the ablation study conducted on the Conformer and Wav2vec2-XLSR architectures are shown in Table 4. Specifically, we conduct ablations on the utilization of a single datastore, two separate datastores, average $n$ neighbor distances for selecting datastore, and scale temperature $t$ to adjust the alternate language distribution. For both models, $k$NN with a single datastore improves the performance compared with the CTC baseline. We can observe that $k$NN with two separate datastores employing the shortest neighbor distances is unstable. While applying the average of $n$ distances, it does improve the robustness of our method and achieve stable improvement on both baselines. Furthermore, the scale temperature $t$ has the most significant impact on the overall MER, adjusting the alternate language probability based on the accuracy of the gated datastore selection mechanism. Employing the gated mechanism with $n$ neighbors’ average distances and scale temperature $t$, we achieve the best performance on both datasets.
(a) Average distance of $n$ neighbors
(b) Scale temperature $t$
Figure 3: Effect of hyper-parameter $n$ and $t$ of Wav2vec2-XLSR on the $DEV$ set.
Table 4: MER (%) of ablation study of Conformer and Wav2vec2-XLSR baselines on $TEST$. indicates that p-value $<$ 0.05 (compared with baseline and baseline with a single datastore).
@@ 4.2 Ablation study
Conformer 26.17 28.82
   +$k$NN (A single datastore) 25.66 27.94
    +$k$NN (Two separate datastores) 25.66 28.04
     +Average $n$ distances 25.57 27.85
      +Scale temperature $t$ 25.02 26.68
Wav2vec2-XLSR 32.65 35.22
   +$k$NN (A single datastore) 32.47 34.94
    +$k$NN (Two separate datastores) 32.35 34.69
     +Average $n$ distances 32.34 34.70
      +Scale temperature $t$ 31.48 33.41
5 Conclusion

### v4

# Font TimesNewRomanPS-BoldMT, TimesNewRomanPSMT is not embedded thanks: Corresponding author. This work has been supported by the National Key R&\&&D Program of China (Grant No.2022ZD0116307) and NSF China (Grant No.62271270)

@@ Abstract
I Introduction
In today’s increasingly interconnected world, the ability to accurately transcribe and understand speech in multilingual and code-switching (CS) environments is of paramount importance [1]. Automatic speech recognition (ASR) systems play a crucial role in facilitating communication across linguistic boundaries. Recent years have witnessed rapid advancements in ASR technology [2, 3]. However, traditional ASR models typically rely on vast amounts of labeled data for each language or dialect, posing a significant challenge in scenarios where such data is scarce or unavailable [4, 5, 6, 7, 8, 9]. Consider, for instance, a scenario where speakers seamlessly transition between two languages, such as English and Chinese, within the same conversation. Traditional ASR systems trained on monolingual datasets struggle to accurately transcribe this CS speech, as similar pronunciations across languages could confuse the model.
@@ I Introduction
Addressing this, researchers have explored zero-shot learning to enable model generalization to CS-ASR task without specific training data. Peng et al. [20] propose a prompt engineering method that enhances the Whisper model [21] for CS-ASR tasks by replacing a single language token in the prompt with two language tokens. Yan et al. [22] suggest a cross-lingual pseudo-labeling modification for monolingual modules to produce transliterations of foreign speech, aiming to circumvent the error propagation of frame-wise language identification (LID) decisions. Despite this advancement, their approach necessitates fine-tuning each monolingual model with cross-lingual pseudo-labeling and using dual encoders during decoding, highlighting ongoing challenges in CS-ASR development.
@@ II-A CS-ASR based on $k$NN-CTC
II-B $k$NN-CTC with gated monolingual datastores
@@
III Experimental setup
III-A Dataset
We utilize ASCEND [29], a Chinese-English dataset for CS-ASR. We split the training set of ASCEND into three parts: Chinese, English, and Mixed (CS-ASR data), denoted as $S_{CN}$, $S_{EN}$, $S_{MIX}$ respectively. Our experiments only employ the Chinese and English subsets for fine-tuning the models and constructing datastores to keep the zero-shot setting. Additionally, we utilize the ASCEND test set (denoted as TEST) and the mixed training set ($S_{MIX}$) as our test sets. The details of each subset are shown in Table I.
TABLE I: Details of splitting the dataset.
@@ III-A Dataset
TABLE II: MER (%) of our proposed method based on the Conformer fine-tuned with $S_{CN}$ and $S_{EN}$.
@@
TABLE III: MER (%) of our proposed method based on the Wav2vec2-XLSR fine-tuned with $S_{CN}$ and $S_{EN}$.
@@
III-B Implementation details
@@
Our experiments are conducted using the open-source toolkit WeNet [30] for Conformer [31] and HuggingFace’s Transformers [32] for Wav2vec2-XLSR [33]. We utilize the open-source checkpoint111 https://github.com/wenet-e2e/wenet/blob/main/docs/pretrained_models.en.md pre-trained by WenetSpeech [34], comprising 12 encoder layers of Conformer and 3 decoder layers of bi-transformer. During finetuning, a learning rate of $5\times 10^{-5}$ and a batch size of 16 with 5000 warmup steps are employed. For Wav2vec2-XLSR baseline, we leverage the open-source checkpoint [^2] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-chinese-zh-cn and fine-tune it with codebase [^3] https://github.com/HLTCHKUST/ASCEND provided by ASCEND [29]. We abbreviate Conformer-CTC to Conformer and Wav2vec2-XLSR-CTC to Wav2vec2-XLSR for simplicity. All reported results are derived using CTC greedy search decoding. It is important to note that our models are fine-tuned exclusively on the subsets $S_{CN}$ and $S_{EN}$, thus preserving the zero-shot CS-ASR setting.
We reimplement the version of $k$NN-CTC (full) [25] using FAISS [35]. For Conformer, we follow $k$NN-CTC to determine the location of keys. For Wav2vec2-XLSR, we select the encoder output as the location of keys. We set $k$ to 1024 following $k$NN-CTC for both baselines. $\lambda$ is approximately 0.25 adjusted using the validation set. Regarding the gated monolingual datastore selection mechanism, we set $n$=300 for the Conformer baseline and $n$=10 for the Wav2vec2-XLSR baseline to compute the average distances of the $n$ nearest neighbors from $D_{CN}$ and $D_{EN}$. We set the distribution calibration temperature $t$ to 5 for the Conformer baseline and 200 for the Wav2vec2-XLSR baseline, adjusting the probability distribution for the alternative language. We adopt the Mixture Error Rate (MER) as outlined in [36] for metrics. MER accounts for both Chinese characters and English words when calculating the edit distance. Errors are tallied separately for Chinese and English based on the language of the reference token.
IV Results
IV-A Zero-shot CS-ASR results
@@
The evaluation results for the Conformer and Wav2vec2-XLSR baselines are summarized in Table II and Table III, respectively. Our approach, employing the $k$NN-CTC with a single datastore $D_{ALL}$, as well as our method utilizing dual monolingual datastores $D_{CN}$ and $D_{EN}$, outperform the fine-tuned CTC method across both baseline models and test sets. Moreover, our integrated method, incorporating dual monolingual datastores with a gated datastore selection mechanism, demonstrates superior performance in all scenarios. This highlights that the utilization of the bilingual datastore introduces distracting noise during the retrieval process, which adversely affects performance. By implementing the gated mechanism to select a monolingual datastore, we effectively mitigate interference from alternate languages, resulting in precise retrieval and improved performance. Our approach achieves a relative MER reduction of 4.4% and 7.4% on the $TEST$ and $S_{MIX}$ sets respectively for the Conformer baseline. Furthermore, for the Wav2vec2-XLSR baseline, we observe a relative MER reduction of 3.6% and 5.1% on the two test sets separately.
@@
To further assess the effectiveness of our proposed method, we compare its performance with Whisper [21] and Prompting Whisper (PW) [20], as shown in Table IV. We reimplement several versions of PW for this comparison. Our method outperforms both Whisper-Small and PW-Small, achieving superior results with fewer parameters and a lower RTF.
TABLE IV: MER (%) of our proposed method based Conformer, in comparison to Whisper and PromptingWhisper (PW) [20].
Model Type # Params $TEST$ RTF
Whisper B 74M 38.79 0.0455
S 244M 27.42 0.0912
PW [20] B 74M 46.38 0.0458
S 244M 25.70 0.0923
Ours (Conformer) - 123M 25.02 0.0151
@@
Figure 2: Visualization of average distances $d_{CN}$ and $d_{EN}$. The green dashed vertical line represents occurrences of CS. The color blue represents Chinese, while orange represents English.
@@
IV-B Ablation study
The results of the ablation study conducted on Wav2vec2-XLSR are shown in Table V. We report Word Error Rate (WER) for English, Character Error Rate (CER) for Chinese and total MER. Specifically, we conduct ablations on the utilization of a single datastore (S1), two separate datastores with a gated datastore selection mechanism (S2), and scale temperature $t$ to adjust the alternate language distribution (S3). Our findings indicate that $k$NN with a single datastore outperforms the CTC baseline. Moreover, $k$NN with two separate datastores further improves performance by reducing decoding interference from extraneous language noise. Furthermore, the scale temperature $t$ has the most significant impact on the overall MER, adjusting the alternate language probability based on the accuracy of the gated datastore selection mechanism. By employing the gated mechanism and scale temperature $t$, we achieve the best performance.
TABLE V: Ablation study of Wav2vec2-XLSR baseline on $TEST$ and $S_{MIX}$.
@@ IV-B Ablation study
CER WER MER CER WER MER
S0 24.30 66.83 32.65 23.98 74.25 35.22
S1 24.23 66.19 32.47 23.65 74.08 34.94
S2 24.04 66.33 32.34 23.49 73.67 34.70
S3 23.62 63.68 31.48 22.53 71.25 33.41
V Conclusion